OpenStudy (anonymous):

Which value is a solution for the equation cot(x/2)=-1? A. 5pi/4 B. 7pi/4 C. 3pi/2 D. 3pi/4

OpenStudy (anonymous):

well you can try solving it by hand. You will have to arccot to both sides cot(x/2)=-1 (x/2)=arccot(-1) x= 2*arccot(-1)

OpenStudy (anonymous):

plug that in to a calculator and see what you get. Or if you dont have a calculator try: http://www.wolframalpha.com/